which equation of the line passes through (3,-2) and is parallel to y=2/3x?
Answer by stanbon(48564)

which equation of the line passes through (3,-2) and is parallel to y=2/3x?
You have y=-2 when x=3 and you have slope = (2/3)
So, -2=(2/3)(3)+b
b=-4
EQUATION:
y=(2/3)x-4
Cheers,
Stan H.